OpenStudy (anonymous):

f(-1)=3, f(2)=6, g(-1)=3, and g(2)=-1 find a). (f-g)(-1) b).(f.g)(-1) c).(f/g)(2) d).(fog)(2)

OpenStudy (anonymous):

a). (f-g)(-1) = 3-3 = 0 b). (f*g)(-1) = 3 * 3 = 9 c). (f/g)(2) = 6/-1 = -6 d). (fog)(2) = g(2) = -1 so f(-1) = 3

OpenStudy (anonymous):

thanks for help

OpenStudy (anonymous):

but don't know how to it